“SUPERCOMBINATION DIFFERENCIES” AS THE METHOD OF DETERMINATION OF A_{1}–A_{2} SPLIT-TINGS IN THE GROUND VIBRATIONAL STATES OF SYMMETRIC TOP MOLECULES: CHD_{3} MOLECULE

\underline{O. N. ULENIKOV}, G. A. ONOPENKO, N. E. TYABAEVA, Tomsk State University, Tomsk, Russia; S. ALANKO, M. KOIVUSAARI, AND R. ANTTILA, University of Oulu, Finland.

“Supercombination differencies” method was derived in order to determine the A_{1}–A_{2} (K=3) ground state splittings in symmetric top molecules from their infrared spectra, and was applied to study of CHD_{3} molecule on the base of high resolution Fourier transform spectra of its 12 absorption bands. Possibility of simultaneous determination of ϵ and h_{3} ambiguous parameters of Watson’s ambiguity theory is discussed. It is shown that this problem can be positively solved on the base of analysis of high resolution rotational structure of the ν_{5} band. New set of ground state energies was determined, which allowed us to explain some earlier unexplained peculiarities in infrared spectra of CHD_{3}.